Using divisibility tests, determine which of the following numbers are divisible by 11: 
70169308 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Sum of digits at odd places = 8 + 3 + 6 + 0 = 17
Sum of digits at even places = 0 + 9 + 1 + 7 = 17
Difference = 17–17 = 0
Since, the difference between sum of digits at odd places and sum of digits at even places is 0.
Hence, 70169308 is divisible by 11

Thanks